Please, I need help with this question not with software. The old fashion way is very helpful/

6. Ryan wants to know if weather affects mood.  To that end, he asks the same 15 people to rate their moods at three different times – when it is raining, when it is sunny and when it is snowing.  The data follows.  Does weather significantly affect mood? SSsubjects = 100

Raining

Sunny

Snowing

M=5, s=1.5, n=15

M=8, s=1.1, n=15

M=8.4, s=3.1, n=15
Answer #1

Ho: weather does not significantly affect mood

H1: weather does significantly affect mood

raining sunny Snowing
count, ni = 15 15 15
mean , x̅ i = 5.000 8.00 8.40
std. dev., si = 1.500 1.100 3.100
sample variances, si^2 = 2.250 1.210 9.610
total sum 75 120 126 321 (grand sum)
grand mean , x̅̅ = Σni*x̅i/Σni =   7.13
square of deviation of sample mean from grand mean,( x̅ - x̅̅)² 4.551 0.751 1.604
TOTAL
SS(between)= SSB = Σn( x̅ - x̅̅)² = 68.267 11.267 24.067 103.6
SS(within ) = SSW = Σ(n-1)s² = 31.500 16.940 134.540 182.980

no. of treatment , k =   3
df between = k-1 =    2
N = Σn =   45
df within = N-k =   42
  
mean square between groups , MSB = SSB/k-1 =    51.8000
  
mean square within groups , MSW = SSW/N-k =    4.3567
  
F-stat = MSB/MSW =    11.8898
P value =   0.0001

anova table
SS df MS F p-value
Between: 103.600 2 51.800 11.890 0.0001
Within: 182.980 42 4.357
Total: 286.580 44

p value <α=0.05, reject Ho

there is enough evidence that weather does significantly affect mood at α=0.05